# Lightweight Dataset Pruning without Full Training via Example Difficulty and Prediction Uncertainty

### v2

@@ Abstract
Recent advances in deep learning rely heavily on massive datasets, leading to substantial storage and training costs. Dataset pruning aims to alleviate this demand by discarding redundant examples. However, many existing methods require training a model with a full dataset over a large number of epochs before being able to prune the dataset, which ironically makes the pruning process more expensive than just training the model on the entire dataset. To overcome this limitation, we introduce a Difficulty and Uncertainty-Aware Lightweight (DUAL) score, which aims to identify important samples from the early training stage by considering both example difficulty and prediction uncertainty. To address a catastrophic accuracy drop at an extreme pruning, we further propose a ratio-adaptive sampling using Beta distribution. Experiments on various datasets and learning scenarios such as image classification with label noise and image corruption, and model architecture generalization demonstrate the superiority of our method over previous state-of-the-art (SOTA) approaches. Specifically, on ImageNet-1k, our method reduces the time cost for pruning to 66% compared to previous methods while achieving a SOTA, specifically 60% test accuracy at a 90% pruning ratio. On CIFAR datasets, the time cost is reduced to just 15% while maintaining SOTA performance.  [^1] Our codebase is available at github.com/dual-pruning.
@@ 1 Introduction
However, many existing pruning methods require training a model with a full dataset over a number of epochs to measure the importance of each sample, which ironically makes the pruning process more expensive than just training the model once on the original large dataset. For instance, several score-based methods (Toneva et al., 2018; Pleiss et al., 2020; Paul et al., 2021; He et al., 2024; Zhang et al., 2024) require training as they utilize the dynamics from the whole training process. Some geometry-based methods, (Xia et al., 2022; Yang et al., 2024) leverage features from the penultimate layer of the trained model, therefore training a model is also required. Hybrid methods (Zheng et al., 2022; Maharana et al., 2023; Tan et al., 2025), which address the difficulty and diversity of samples simultaneously, still hold the same limitation as they use existing score metrics. Having to compute the dot product of learned features to get the neighborhood information makes them even more expensive to utilize.
@@
To address this issue, we introduce Difficulty and Uncertainty-Aware Lightweight (DUAL) score, a metric that measures the importance of samples in the early stage of training by considering both prediction uncertainty and the example difficulty. Additionally, at the high pruning ratio—when the selected subset is scarce—we propose pruning-ratio-adaptive Beta sampling, which intentionally includes easier samples with lower scores to achieve a better representation of the data distribution (Sorscher et al., 2022; Zheng et al., 2022; Acharya et al., 2024).
Experiments conducted on CIFAR and ImageNet datasets under various learning scenarios verify the superiority of our method over previous SOTA methods. Specifically, on ImageNet-1k, our method reduces the time cost to 66% compared to previous methods while achieving a SOTA 60% test accuracy at the pruning ratio of 90%. On the CIFAR datasets, as illustrated in Figure 1, our method reduces the time cost to just 15% while maintaining SOTA performance. Especially, our method shows a notable performance when artificial noise is added.
@@ 2 Related Works
Data pruning aims to remove redundant examples, keeping the most informative subset of samples, namely the coreset. Research in this area can be broadly categorized into two groups: score-based and geometry-based methods. Score-based methods define metrics representing the difficulty or importance of data points to prioritize samples with high scores. Geometry-based methods, whereas, focus more on keeping a good representation of the true data distribution. Recent studies proposed hybrid methods that incorporate the example difficulty score with the diversity of coreset.
Score-based. EL2N (Paul et al., 2021) calculates $L2$ norms of the error vector as an approximation of the gradient norm. Entropy (Coleman et al., 2020) quantifies the information contained in the predicted probabilities at the end of training. However, the outcomes of such “snapshot” methods differ significantly from run to run, making it difficult to obtain a reliable score in a single run, as can be seen in Figure 7, Appendix B.
@@
Methods using training dynamics offer more reliability as they incorporate information throughout an entire run of training. Forgetting (Toneva et al., 2018) score counts the number of forgetting events, a correct prediction on a data point is flipped to a wrong prediction during training. AUM (Pleiss et al., 2020) accumulates the gap between the target probability and the second-highest prediction probability. Dyn-Unc (He et al., 2024), which strongly inspired our approach, prioritizes the uncertain samples rather than typical easy samples or hard ones during model training. The prediction uncertainty is measured by the variation of predictions in a sliding window, and the score averages the variation throughout the whole training process. TDDS (Zhang et al., 2024) averages differences of Kullback-Leibler divergence loss of non-target probabilities for $T$ training epochs, where $T$ is highly dependent on the pruning ratio. Taking the training dynamics into account proves useful for pruning because it allows one to differentiate informative but hard samples from ones with label noise (He et al., 2024). However, despite the stability and effectiveness of these methods, they fail to provide cost-effectiveness as it requires training the model on the entire dataset.
@@
Geometry-based. Geometry-based methods focus on reducing redundancy among selected samples to provide better representation. SSP (Sorscher et al., 2022) selects the samples most distant from k-means cluster centers, while Moderate (Xia et al., 2022) focuses on samples with scores near the median. However, these methods often compromise generalization performance as they underestimate the effectiveness of difficult examples.
@@
Recently, hybrid approaches have emerged that harmonize both difficulty and diversity. CCS (Zheng et al., 2022) partitions difficulty scores into bins and selects an equal number of samples from each bin to ensure a balanced representation. $\mathbb{D}^{2}$ (Maharana et al., 2023) employs a message-passing mechanism with a graph structure where nodes represent difficulty scores and edges encode neighboring representations, facilitating effective sample selection. BOSS (Acharya et al., 2024) introduces a Beta function for sampling based on difficulty scores, which resembles our pruning ratio-adaptive sampling; we discuss the key difference in Section 3.3. Our DUAL pruning is a score-based approach as it considers difficulty and uncertainty by the score metric. Additionally, diversity is introduced through our proposed Beta sampling, making it a hybrid approach.
@@ 3.1 Preliminaries
where ¯$\bar{{\mathbb{P}}}_{k}:=\frac{\sum_{j=0}^{J-1}{\mathbb{P}}_{k+j}(y\mid{\bm{x}})}{J}$ is the average prediction of the model over the window $[k,k+J-1]$. Then taking the average of the uncertainty throughout the whole training process leads to the Dyn-Unc score:
@@ 3.2 Difficulty & Uncertainty-Aware Lightweight Score
Following the approach of Swayamdipta et al. (2020) and He et al. (2024), we analyze data points from ImageNet-1k based on the mean and standard deviation of predictions during training, as shown in Figure 2. We observe data points typically “flow” along the “moon” from bottom to top direction. Data points starting from the bottom-left region with a low prediction mean and low standard deviation move to the middle region with increased mean and standard deviation, and those starting in the middle region drift toward the upper-left region with a high prediction mean and smaller standard deviation. This phenomenon is closely aligned with existing observations that neural networks typically learn easy samples first, then treat harder samples later  (Bengio et al., 2009; Arpit et al., 2017; Jiang et al., 2020; Shen et al., 2022). In other words, we see that the uncertainty of easy samples rises first, and then more difficult samples start to move and show an increased uncertainty score.
Figure 2 further gives a justification for this intuition. In Figure 2(a), samples with the highest Dyn-Unc scores calculated at epoch 60 move upward by the end of training at epoch 90. It means that if we measure the Dyn-Unc score at the early stage of training, it gives the highest scores to relatively easy samples rather than the most informative samples. It seems undesirable that it results in poor test accuracy on its coreset as shown in Figure 8 of Appendix B.
To capture the most useful samples that are likely to contribute significantly to dynamic uncertainty during the whole training process (of 90 epochs) at the earlier training stage ($e.g.$ epoch of 60), we need to target the samples located near the bottom-right region of the moon-shaped distribution, as Figure 2(b) illustrates. Inspired by this observation, we propose a scoring metric that identifies such samples by taking the uncertainty of the predictions and the prediction probability into consideration.
(a) Dyn-Unc score calculated at epoch $T=60$.
(b) Dyn-Unc score calculated at epoch $T=90$.
Figure 2: The left column visualizes the prediction mean and standard deviation for each data point collected up to epoch 60, while the right column stands for epoch 90. Samples are colored by normalized Dyn-Unc score for each row. The expected cases of the original Dyn-Unc, where the score computation time step is the same as the final epoch of training, are marked with bold outlines.
@@
The DUAL score reflects training dynamics by leveraging prediction probability across several epochs It provides a reliable estimation to identify the most uncertain examples.
A theoretical analysis of a toy example further verifies the intuition above. Consider a linearly separable binary classification task $\left\{({\bm{x}}_{i}\in{\mathbb{R}}^{n},y_{i}\in\{\pm 1\})\right\}_{i=1}^{N}$, where $N=2$ with $\lVert{\bm{x}}_{1}\rVert^{2}$ $\ll$ $\langle{\bm{x}}_{1},{\bm{x}}_{2}\rangle<\lVert{\bm{x}}_{2}\rVert^{2}$. Without loss of generality, we set $y_{1}=y_{2}=+1$. A linear classifier, $f({\bm{x}};{\bm{w}})={\bm{w}}^{\top}{\bm{x}}$, is employed as the model in our analysis. The parameter ${\bm{w}}$ is initialized at zero and updated by gradient descent. Soudry et al. (2018) prove that the parameter of linear classifiers diverges to infinity, but directionally converges to the $L_{2}$ maximum margin separator. This separator is determined by the support vectors closest to the decision boundary. If a valid pruning method encounters this task, then it should retain the point closer to the decision boundary, which is ${\bm{x}}_{1}$ in our case, and prune ${\bm{x}}_{2}$.
Due to its large norm, ${\bm{x}}_{2}$ exhibits higher score values in the early training stage for both uncertainty and DUAL scores. It takes some time for the model to predict ${\bm{x}}_{1}$ with high confidence, which increases its uncertainty level and prediction mean, as well as for scores of ${\bm{x}}_{1}$ to become larger than ${\bm{x}}_{2}$ as training proceeds. In Theorem 3.1, we show through a rigorous analysis that the moment of such a flip in order happens strictly earlier for DUAL than uncertainty.
@@
Define $\sigma(z)\coloneq(1+e^{-z})^{-1}$. Let $V_{t;J}^{(i)}$ be the variance and $\mu_{t;J}^{(i)}$ be the mean of $\sigma(f({\bm{x}}_{i};{\bm{w}}_{t}))$ within a window from time $t$ to $t+J$. Denote $T_{v}$ and $T_{vm}$ as the first time step when $V_{t;J}^{(1)}>V_{t;J}^{(2)}$ and $V_{t;J}^{(1)}(1-\mu_{t;J}^{(i)})>V_{t;J}^{(2)}(1-\mu_{t;J}^{(2)})$ occur, respectively. If the learning rate is small enough, then $T_{vm}<T_{v}$.
@@
Figure 3: DUAL score also targets uncertain samples during the early epoch. In the end, selected samples are finally located in the most uncertain region.
Empirically, as shown in Figure 3, the DUAL score targets data points in the bottom-right region during the early training phase, which eventually evolve to the middle-rightmost part by the end of training. This verifies that DUAL pruning identifies the most uncertain region faster than Dyn-Unc both in theory and practice. The distinction arises from the additional consideration of an example difficulty in our method. We believe that this adjustment leads to improved generalization performance compared to Dyn-Unc, as verified through various experiments in later sections.
However, score-based methods, including our method, suffer from limitations due to biased representations, leading to poor coreset test accuracy at high pruning ratios. ratios. To address this, we propose an additional strategy to adaptively select samples according to a ratio for pruning.
3.3 Pruning Ratio-Adaptive Sampling
Since the distribution of difficulty scores is dense in high-score samples, selecting only the highest-score samples may result in a biased model (Zhou et al., 2023; Maharana et al., 2023; Choi et al., 2024). To address this, we design a sampling method to determine the subset ${\mathcal{S}}\subset{\mathcal{D}}$, rather than simply pruning the samples with the lowest score. We introduce a Beta distribution that varies with the pruning ratio. The primary objective of this method is to ensure that the selected subsets gradually include more easy samples into the coreset as the pruning ratio increases.
However, the concepts of “easy” and “hard” cannot be distinguished solely based on uncertainty, or DUAL score. To address this, we use the prediction mean again for sampling. We utilize the Beta probability density function (PDF) to define the selection probability of each sample. First, we assign each data point a corresponding PDF value based on its prediction mean and weight this probability using the DUAL score. The weighted probability with the DUAL score is then normalized so that the sum equals 1, and then used as the sampling probability. To be clear, sampling probability is for selecting samples, not for pruning. Therefore, for each pruning ratio $r$, we randomly select $(1-r)\cdot n$ samples without replacement, where sampling probabilities are given according to the prediction mean and DUAL score as described. The detailed algorithm for our proposed pruning method with Beta sampling is provided in Algorithm 1, Appendix C.
We design the Beta PDF to assign a sampling probability concerning a prediction mean as follows:
where $C>0$ is a fixed constant, and the $\mu_{D}$ stands for the prediction mean of the highest score sample. Recalling that the mean of Beta distribution is $\frac{\alpha_{r}}{\alpha_{r}+\beta_{r}}$, the above choice makes the mean of Beta distribution move progressively with $r$, starting from $\mu_{\mathcal{D}}$ ($r\simeq 0$, small pruning ratio) to one. In other words, with growing $r$, this Beta distribution becomes skewed towards the easier region ($r\rightarrow 1$, large pruning ratio), which in turn gives more weight to easy samples. The tendency of evolving should be different with datasets, thus a hyperparameter $c_{\mathcal{D}}\geq 1$ is used to control the rate of evolution of the Beta distribution. Specifically, the choice of $c_{\mathcal{D}}$ depends on the complexity of the initial dataset. For smaller and more complex datasets, setting $c_{\mathcal{D}}$ to a smaller value retains more easy samples. For larger and simpler datasets, setting $c_{\mathcal{D}}$ to a larger value allows more uncertain samples to be selected. (For your intuitive understanding, please refer to Figure 16 and Figure 17 in the Appendix C.) This is also aligned with the previous findings from Sorscher et al. (2022); if the initial dataset is small, the coreset is more effective when it contains easier samples, while for a relatively large initial dataset, including harder samples can improve generalization performance.
Remark.
4 Experiments
Table 1: Comparison of test accuracy between the DUAL score method and existing techniques using ResNet-18 on CIFAR-10 and CIFAR-100 datasets. Training the model on the full dataset achieves an average test accuracy of 95.30% on CIFAR-10 and 78.91% on CIFAR-100. The best result in each pruning ratio is highlighted in bold.
@@ 4 Experiments
chat Pruning Rate ($\rightarrow$) 30% 50% 70% 80% 90% 30% 50% 70% 80% 90%

### v3

@@ Abstract
Recent advances in deep learning rely heavily on massive datasets, leading to substantial storage and training costs. Dataset pruning aims to alleviate this demand by discarding redundant examples. However, many existing methods require training a model with a full dataset over a large number of epochs before being able to prune the dataset, which ironically makes the pruning process more expensive than just training the model on the entire dataset. To overcome this limitation, we introduce the Difficulty and Uncertainty-Aware Lightweight (DUAL) score, which aims to identify important samples from the early training stage by considering both example difficulty and prediction uncertainty. To address a catastrophic accuracy drop at an extreme pruning ratio, we further propose a pruning ratio-adaptive sampling using Beta distribution. Experiments on various datasets and learning scenarios such as image classification with label noise and image corruption, and model architecture generalization demonstrate the superiority of our method over previous state-of-the-art (SOTA) approaches. Specifically, on ImageNet-1k, our method reduces the time cost for pruning to 66% compared to previous methods while achieving a SOTA 60% test accuracy at a 90% pruning ratio. On CIFAR datasets, the time cost is reduced to just 15% while maintaining SOTA performance. Implementation is available at github/dual-pruning.
Machine Learning, ICML
@@ 1 Introduction
However, many existing pruning methods require training a model with a full dataset over a number of epochs to measure the importance of each sample, which ironically makes the pruning process more expensive than just training the model once on the original large dataset. For instance, several score-based methods (Toneva et al., 2018; Pleiss et al., 2020; Paul et al., 2021; He et al., 2024; Zhang et al., 2024) require training as they utilize the dynamics from the whole training process. Some geometry-based methods (Xia et al., 2022; Yang et al., 2024) leverage features from the penultimate layer of the trained model, therefore training a model is also required. Hybrid methods (Zheng et al., 2022; Maharana et al., 2023; Tan et al., 2025), which address the difficulty and diversity of samples simultaneously, still hold the same limitation as they use existing score metrics. Having to compute the dot product of learned features to get the neighborhood information makes them become even more expensive to utilize.
@@
In order to address this issue, we introduce the Difficulty and Uncertainty-Aware Lightweight (DUAL) score, which can measure the importance of samples in the early stage of the training process by considering example difficulty and the prediction uncertainty. Additionally, for the high pruning ratio—when the selected subset is scarce—we propose pruning-ratio-adaptive Beta sampling, to intentionally include easier samples which have lower scores to achieve a better representation of the data distribution (Sorscher et al., 2022; Zheng et al., 2022; Acharya et al., 2024).
Experiments conducted on CIFAR and ImageNet datasets under various learning scenarios verify the superiority of our method. Specifically, on ImageNet-1k, our method reduces the time cost to 66% compared to previous methods while achieving a SOTA performance, 60% test accuracy at the pruning ratio of 90%. On the CIFAR datasets, as illustrated in Figure 1, our method reduces the time cost to just 15% while maintaining SOTA performance. Especially, our proposed method shows a notable performance when the dataset contains noise.
@@ 2 Related Works
Data pruning aims to remove redundant examples, keeping the most informative subset of dataset, namely the coreset. Research in this area can be broadly categorized into two groups: score-based and geometry-based methods. Score-based methods define metrics to measure the importance of data points, then prioritize high-scoring samples. On the other hand, geometry-based methods focus on presenting a better representation of the true data distribution. Recent studies propose hybrid methods which incorporate the example difficulty score with the diversity of the coreset.
Score-based.
@@
EL2N (Paul et al., 2021) calculates L2 norms of the error vector, as an approximation of the gradient norm. Entropy (Coleman et al., 2020) quantifies the information contained in the predicted probabilities at the end of training. However, the outcomes of such “snapshot” methods need multiple runs to be stabilized, as shown in Figure 7, Appendix B. AUM (Pleiss et al., 2020) accumulates the gap between the target probability and the second-highest prediction probability. Methods that utilize training dynamics throughout the entire training offer more stable measures. Forgetting (Toneva et al., 2018) score counts the number of forgetting events, where a correct prediction is flipped to a wrong prediction during training process. Dyn-Unc (He et al., 2024), which strongly inspired our approach, prioritizes the most uncertain samples rather than typical easy or hard samples during model training. The uncertainty is measured by the variation of predictions in a sliding window, and the score averages the variation throughout the whole training process. TDDS (Zhang et al., 2024) averages differences of Kullback-Leibler divergence loss of non-target probabilities for $T$ training epochs, where $T$ is highly dependent on the pruning ratio. The information from training dynamics proves useful for pruning because it allows one to differentiate hard but useful samples from noisy ones (He et al., 2024). However, despite its stability and effectiveness, previous methods fail to guarantee cost-effectiveness as they require at least one full training of the model on the entire dataset.
Geometry-based.
@@
Geometry-based methods concentrate on providing a better representation by minimizing the redundancy of selected samples. SSP (Sorscher et al., 2022) selects the samples most distant from k-means cluster centers, while Moderate (Xia et al., 2022) prefers samples near the median. However, these methods often compromise generalization performance, since they underestimate the effect of difficult examples.
@@
Recently, hybrid approaches have emerged that harmonize both difficulty and diversity. CCS (Zheng et al., 2022) partitions difficulty scores into bins and selects an equal number of samples from each bin to ensure balanced representation. $\mathbb{D}^{2}$ (Maharana et al., 2023) employs a message-passing mechanism with a graph structure where nodes represent difficulty scores and edges encode neighboring representations, facilitating effective sample selection. BOSS (Acharya et al., 2024) introduces a Beta function for importance sampling based on difficulty scores, which resembles our pruning ratio-adaptive sampling; we discuss the differences in Section 3.3. SIMS (Grosz et al., 2024) defines SIM score using class separability, data integrity, and model uncertainty, and then integrates sampling strategy. Our DUAL pruning defines a score metric with difficulty and uncertainty, and it becomes a hybrid approach when the score is combined with our proposed Beta sampling.
@@ 3.1 Preliminaries
where ¯$\bar{{\mathbb{P}}}_{k}:=\frac{\sum_{j=0}^{J-1}{\mathbb{P}}_{k+j}(y\mid{\bm{x}})}{J}$ is the average prediction of the model over the window $[k,k+J-1]$. Then taking the average of the uncertainty throughout the whole training process leads to Dyn-Unc score:
@@ 3.2 Difficulty & Uncertainty-Aware Lightweight Score
Following the approach of Swayamdipta et al. (2020) and He et al. (2024), we analyze data points from ImageNet-1k based on the mean and standard deviation of predictions during training, as shown in Figure 2. We observe data points typically flow along the moon from bottom to top. Data points starting from the bottom left region with low prediction mean and low standard deviation move to the middle region with increased mean and standard deviation, and those starting at the middle region drift toward to the upper left region with high prediction mean and smaller standard deviation. This phenomenon is closely aligned with existing observations that neural networks typically learn easy samples first, then treat harder samples later (Bengio et al., 2009; Arpit et al., 2017; Jiang et al., 2020; Shen et al., 2022). In other words, we see that the uncertainty of easy samples rises first, and then more difficult samples start to move and show increased uncertainty score.
Figure 2 further gives a justification for this intuition. In Figure 2(a), samples with the highest Dyn-Unc scores calculated at epoch 60 move upward by the end of training at epoch 90. This means that if we measure Dyn-Unc score at the early stage of training, it gives the highest scores to relatively easy samples rather than the most informative samples. It seems undesirable that it results in poor test accuracy on its coreset, as shown in Figure 8 in Appendix B.
To capture the most useful samples that are likely to contribute significantly to Dyn-Unc during the whole training process (of 90 epochs) at the earlier training stage ($e.g.$ epoch of 60), we need to target the samples located near the bottom-right region of the moon-shaped distribution, as Figure 2(b) illustrates. Inspired by this observation, we design a scoring metric that identifies such samples by taking both the uncertainty of the predictions and the prediction probability into consideration.
(a) Score calculated at epoch $T=60$.
(b) Score calculated at epoch $T=90$.
Figure 2: The left column (“Epoch 60”) shows the prediction mean and standard deviation, computed using the predicted target probabilities up to epoch 60. The right column (“Epoch 90”) shows corresponding values up to epoch 90. In each row, samples are colored by the normalized Dyn-Unc score computed at epoch 60 for Figure 2(a) and at epoch 90 for Figure 2(b). The epoch at which the score was computed is indicated by a bold outline for each row.
@@
The DUAL score reflects training dynamics by leveraging prediction probability across several epochs, providing a reliable estimation to identify the most uncertain examples.
A theoretical analysis of a toy example further verifies the intuition above. Consider a linearly separable binary classification task $\left\{({\bm{x}}_{i}\in{\mathbb{R}}^{n},y_{i}\in\{\pm 1\})\right\}_{i=1}^{N}$, where $N=2$ with $\lVert{\bm{x}}_{1}\rVert$ $\ll$ $\langle{\bm{x}}_{1},{\bm{x}}_{2}\rangle<\lVert{\bm{x}}_{2}\rVert$. Without loss of generality, we set $y_{1}=y_{2}=+1$. A linear classifier, $f({\bm{x}};{\bm{w}})={\bm{w}}^{\top}{\bm{x}}$, is employed as the model in our analysis. The parameter ${\bm{w}}$ is initialized at zero and updated by gradient descent. Soudry et al. (2018) prove that the parameter of linear classifiers diverges to infinity, but directionally converges to the $L_{2}$ maximum margin separator. This separator is determined by the support vectors closest to the decision boundary. If a valid pruning method encounters this task, then it should retain the point closer to the decision boundary, which is ${\bm{x}}_{1}$ in our case, and prune ${\bm{x}}_{2}$. Due to its large norm, ${\bm{x}}_{2}$ exhibits higher score values in the early training stage, for both Dyn-Unc and DUAL scores. It takes some time for the model to make prediction on ${\bm{x}}_{1}$ with large confidence to increase its uncertainty level as well as prediction mean, and the scores for ${\bm{x}}_{1}$ eventually become larger than those for ${\bm{x}}_{2}$ as training proceeds. In Theorem 3.1, we show through a rigorous analysis that the moment of such a flip in order happens strictly earlier for DUAL than for uncertainty.
@@
Define $\sigma(z)\coloneq(1+e^{-z})^{-1}$. Let $S_{t;J}^{(i)}$ be the standard deviation and $\mu_{t;J}^{(i)}$ be the mean of $\sigma(f({\bm{x}}_{i};{\bm{w}}_{t}))$ within a window from time $t$ to $t+J-1$. Denote $T_{v}$ and $T_{vm}$ as the first time when $S_{t;J}^{(1)}>S_{t;J}^{(2)}$ and $S_{t;J}^{(1)}(1-\mu_{t;J}^{(i)})>S_{t;J}^{(2)}(1-\mu_{t;J}^{(2)})$ occurs, respectively. If the learning rate is small enough, then $T_{vm}<T_{v}$.
@@
Empirically, as shown in Figure 3, the DUAL score targets data points in the bottom-right region during the early training phase, which eventually evolve to the middle-rightmost part by the end of training. This verifies that DUAL pruning identifies the most uncertain region faster than Dyn-Unc both in theory and practice. The differences arise from the additional consideration of example difficulty in our method. We believe this adjustment leads to improved generalization performance compared to Dyn-Unc, as verified through various experiments in later sections.
Figure 3: Our DUAL score targets similar uncertain samples in the early epoch of 60 (highlighted in bold). Selected samples are finally located in the most uncertain region when the whole training processes are considered.
However, score-based approaches including our method, suffer from accuracy drop at the high pruning ratio due to biased representations. To address this, we propose an additional sampling strategy that adaptively selects samples regarding the coreset size.
Table 1: Comparison of test accuracy between the DUAL score method and existing coreset selection techniques using ResNet-18 on CIFAR-10 and CIFAR-100 datasets. Training the model on the full dataset achieves an average test accuracy of 95.30% on CIFAR-10 and 78.91% on CIFAR-100. The best result in each pruning ratio is highlighted in bold.
@@
Pruning Rate ($\rightarrow$) 30% 50% 70% 80% 90% 30% 50% 70% 80% 90%
@@
3.3 Pruning Ratio-Adaptive Sampling
Since the distribution of difficulty scores is dense in high-score samples, selecting only the highest-score samples may result in a biased model (Zhou et al., 2023; Maharana et al., 2023; Choi et al., 2024). To address this, we design a sampling method to determine the subset ${\mathcal{S}}\subset{\mathcal{D}}$, rather than simply pruning the samples with the lowest scores. We introduce a Beta distribution that varies with the pruning ratio. The primary objective of this method is to ensure that the selected subsets gradually include more easy samples into the coreset as the pruning ratio increases.
However, the concepts of “easy” and “hard” cannot be distinguished solely based on uncertainty or DUAL score. To address this, we use the prediction mean again for sampling. We utilize the Beta probability density function (PDF) to define the selection probability of each sample. First, we assign each data point a corresponding PDF value based on its prediction mean and weight this probability using the DUAL score. The weighted probability with the DUAL score is then normalized to the range $[0,1]$, then used as the sampling probability. We clarify that sampling probability is for selecting samples, not for pruning. Therefore, for each pruning ratio $r$, we randomly select $(1-r)\cdot n$ samples without replacement, where sampling probabilities are given according to the prediction mean and DUAL score as described. The detailed algorithm for our proposed pruning method is provided in Algorithm 1, Appendix C.
We design the Beta PDF function to assign a sampling probability concerning a prediction mean as follows:
where $C>0$ is a fixed constant, and the $\mu_{D}$ stands for the prediction mean of the highest score sample. Recalling that the mean of Beta distribution is $\frac{\alpha_{r}}{\alpha_{r}+\beta_{r}}$, the above choice makes the mean of Beta distribution moves progressively with $r$, starting from $\mu_{\mathcal{D}}$ ($r\simeq 0$, small pruning ratio) to one. In other words, with growing $r$, this Beta distribution becomes skewed towards the easier region ($r\rightarrow 1$, large pruning ratio), which in turn gives more weight to easy samples.
The tendency of evolving should be different with datasets, thus a hyperparameter $c_{\mathcal{D}}\geq 1$ is used to control the rate of evolution of the Beta distribution. Specifically, the choice of $c_{\mathcal{D}}$ depends on the complexity of the initial dataset. For smaller and more complex datasets, setting $c_{\mathcal{D}}$ to a smaller value retains more easy samples. For larger and simpler datasets, setting $c_{\mathcal{D}}$ to a larger value allows more uncertain samples to be selected. (For your intuitive understanding, please refer to Figure 20 and Figure 21 in the Appendix C.) This is also aligned with the previous findings from Sorscher et al. (2022); if the initial dataset is small, the coreset is more effective when it contains easier samples, while for a relatively large initial dataset, including harder samples can improve generalization performance. More descriptions for our Beta sampling are provided in Appendix C.
Remark.
Remark.
SIMS (Grosz et al., 2024) also proposes a ratio-adaptive sampling strategy, applying importance weights over the original score distribution. However, it assumes a normal distribution of scores, which does not hold in practice (see Figure 2 of (Grosz et al., 2024)). In contrast, our sampling method, by not relying on any specific score distribution, remains robust across diverse datasets.
4 Experiments